### v1

@@ 2.2 Societal and Human-Centric Values
Beyond economic and business values, societal and human-centric values, which cover other important aspects, are also crucial for businesses and platforms.
Societal and human-centric values for stakeholders in recommender systems focus on ensuring that these systems operate in ways that prioritize humans individually and society as a whole. We find that there are 4 themes of societal and human-centric values for stakeholders in recommender systems that are relevant in the light of evaluation: (i) usefulness, (ii) well-being, (iii) legal and human rights, and (iv) public discourse and safety [69, 70].
Usefulness and enjoyment means that recommendations should meet the needs and expectations of its stakeholders effectively and efficiently [35]. For example, in the case of a music recommender system, users should be able, via the recommender system, to discover new music that they might enjoy and match their tastes. At the same time, usefulness refers to the recommender system’s ability to help music artists get their outputs recommended to potentially interested listeners. Control and privacy is a closely related value that pertains to the degree of influence and customization stakeholders might have over the recommendations that are generated. This includes privacy aspects in a way that users might want to control the amount of their (music) preference data that is shared with the recommender system [52, 69].
Well-being refers to the recommender system’s ability to help its stakeholders to feel satisfied. In the case of a music recommender system, this means that recommendations should influence the experience with the music streaming platform positively, e.g., provide music recommendations to help listeners relax or relieve stress [34]. In this respect, well-being is related to emotional, mental, and physical health. Other related values are connection, community and social bonding, e.g., to enable users to connect with like-minded music listeners or to enable music artists to contribute their outputs to a specific community. Thus, also reputation, recognition and acknowledgement might be valuable for some stakeholders, e.g., to support music artists in getting their contributions recognized by music listeners [49]. Personal growth and development might also be values contributing to well-being in the sense that, e.g., music recommendations could help people explore new music styles and genres, supporting exploration and self-discovery [6].
@@
Fairness can be related to diversity when the goal is to ensure, for example, that diverse articles and styles are represented in recommendation outputs. Diversity may also have listener-oriented benefits, e.g., help music listeners explore artists that might be new to them [58, 13]. A recommender system might enable freedom of expression as well as accessibility and inclusiveness by allowing, e.g., music artists to promote their content independent of the genre or popularity of their music [3, 59]. At the same time, recommender systems should enable users to access the content that they like and enjoy, even when their taste does not match the one of the majority of other music listeners [20, 38]. Transparency and trustworthiness might also be an important value for all stakeholders of a recommender system. For instance, music artists might be interested in why they are ranked at a specific position and music listeners might be interested in why a specific artist was recommended to them [65].
Values in the area of public discourse and safety are related to a multitude of societal and human-centric aspects. Here, societal benefit goes beyond the satisfaction of individual stakeholders. As an example, a music streaming platform might be interested in fostering cultural enrichment by the recommendation of a diverse set of music [73]. This is related to the value of tradition and history, for instance, by recommending local and traditional music, which might be hard to find without the recommender system [21, 42]. The environmental sustainability might also be an important value for some recommender systems stakeholders. This may involve implementing energy-efficient recommendation models within the platforms, or promoting local music artists whose concerts offer the opportunity for attendance without requiring extensive travel [46]. Finally, safety is concerned with users not being exposed to recommendations of disturbing ethically questionable, or age-inappropriate content. In the case of music recommendations, this could refer to sexist or racist music tracks [47, 56].
@@ 2.3 Values in Practice
As we mentioned earlier, the concept of value can be perceived as abstract. Nevertheless, in the context of evaluation of multistakeholder recommender systems, we must be able to somehow quantify it, if the aim is to determine ‘goodness’ for all involved. This task we turn to in Section 3.
@@ 3 Methodology
As we previously noted, evaluating recommender systems is a contextually situated problem: different domains, recommendation tasks, and contexts require specific metrics and evaluation setups tailored to that specific recommendation scenario. Multistakeholder evaluation, where the perspectives of other stakeholders are taken into account in addition to that of the consumer, only increases the potential complexity of the evaluation. The complexity of multistakeholder evaluation is demonstrated by the richness and variety of the examples described in Section 4. As a result of this complexity, prescribing exactly which methods to use in which order is impractical. Instead, we attempt to describe best meta-practices for conducting successful multistakeholder evaluation in this section, divided into different stages. We consider this process to be iterative, as findings in a later stage can necessitate returning to an earlier stage, for instance, when learning of a new relevant stakeholder to include or when value shifts occur in one or more stakeholders.
Recall that, in our discussion, we assume that we seek to evaluate an existing recommender system, one that has already been developed to provide a particular recommendation function. Of course, planning for a system’s evaluation should be part of its development and stakeholder consultation should be prioritized in the design and implementation of a multistakeholder recommender system.
@@ 3.1 Stakeholders
Consumers (or users) have historically played (and continue to play) a central role in recommender systems evaluation. As a result, a common next step would be profiling the consumer stakeholder and the different subgroups this stakeholder category may represent. In addition to interviews with the system stakeholders, any existing market or user research on the user base of the recommendation platform could serve as a valuable foundation for identifying representative subgroups within this user base. A literature review aimed at identifying similar or related recommendation scenarios could also help identify different user groups, especially groups that may be underrepresented in the market research for whatever reason. The system stakeholder should be able to facilitate access to these subgroups, for instance through user research panels, surveys on the website, or customer mailing lists. It is important to recruit a diverse and representative sample of consumers to represent the customer stakeholder and ensure all voices are heard in the evaluation process. Customers should be interviewed or surveyed about what values matter to them in this recommendation scenario (and their relative importance), what goals they have, and how and when they envision using the recommender system. If representative, the principle of saturation could be useful in guiding the sample size required: if additional participants do not reveal any new values, goals, or usage scenarios, then the sample should be representative of the customer stakeholder. Consumers are also a valuable source for identifying possible downstream stakeholders that are worth including in the evaluation process.
@@ 3.3 Evaluation Metrics
Formal evaluation metrics provide a way to measure the extent to which the goals of various stakeholders are achieved, i.e., they are measurable proxies towards goals. For example, both consumers and providers are likely to be interested in recommendation accuracy, consumers may be further interested in item discoverability (diversity, novelty, coverage), providers are likely interested in increasing revenue and engagement, and the third-party stakeholders (for instance, regulators) are likely to be interested in consumer-protection-related metrics (representation, fairness, etc.).
Multiple metrics can measure the success of the same goal, depending on the point of view or the aspect we want to highlight. For example, there are different metrics to measure accuracy, e.g., Normalized Discounted Cumulative Gain (NDCG), Mean Reciprocal Rank (MRR), or Recall; we may measure the overall number of items sold in a specific period or a specific geographical area, the items from the long-tail and the short-head, etc. Depending on the goal, we may have metrics not targeting the overall population of users and stakeholders available in the system.
@@
3.4 Strategies for Overall Multistakeholder Evaluation
Identifying the list of important evaluation metrics $(m_{1},\ldots,m_{n})$, as discussed above, provides the ability to evaluate (i.e., to score) a given recommender system $R$ in a multidimensional manner. More formally, $\mathbf{S}(R)=(s_{1},\ldots,s_{n})$, where $s_{i}$ is the performance of $R$ with respect to measure $m_{i}$, i.e., $s_{i}=m_{i}(R)$. Having multiple evaluation measures raises an important challenge of how to determine the overall (i.e., multistakeholder, multiobjective) performance of the system [76]. In particular, given two candidate recommender systems $R_{A}$ and $R_{B}$, where each of which can be evaluated according to the stated list of metrics, $\mathbf{S}(R_{A})$ and $\mathbf{S}(R_{B})$, how to design a multistakeholder/multiobjective evaluation mechanism $\prec_{M}$ that allows to determine whether system $R_{B}$ has superior overall performance to system $R_{A}$, i.e., $\mathbf{S}(R_{A})\prec_{M}\mathbf{S}(R_{B})$?
Example strategies for developing multistakeholder/multiobjective evaluation mechanisms $\prec_{M}$ include:
@@ 3.5 Practical guidelines
  • Identification of values and goals. The goals of the recommender system and expectations for what makes a recommendation good may vary considerably by stakeholder and context. Each stakeholder has different goals and expectations for the recommender system, and these are directly or indirectly tied to the values that matter to them. Qualitative research methods are particularly useful for identifying these values and goals. It is important to keep domain differences in mind when identifying the values and goals. In some domains, certain stakeholders might be affected more seriously by the selection of recommendations provided by the system. It is therefore recommended to undertake a risk analysis at this stage to properly understand the extent of these risks.
  • Selection of evaluation metrics. Once values and goals have been mapped, they have to be represented by measurable entities, e.g., by selecting existing metrics from prior research that are relevant for a given application context by designing new ones. Part of this process includes determining which metrics to prioritize in the evaluation of the system, as typically multiple metrics can measure the success of the same goal depending on which perspective to highlight. These metrics can vary considerably in terms of being more qualitative or quantitative in nature, or in terms of representing more short-term or long-term interests. All selected metrics must be validated by the target stakeholders to check whether they are representative of their goals.
@@ 4 Example Applications and Metrics
Deriving an evaluation metric requires working from a construct, an abstract quality of the recommendation process that we would like to understand, to a concrete proxy of that construct that can be measured and designing a methodology to measure it. The application-specificity of multistakeholder evaluation means that it is difficult to provide such analysis in a general way. With that in mind, we present several specific examples, which serve as means to guide how researchers and industry practitioners might proceed when developing such metrics.
In each of these hypothetical examples, we select a particular stakeholder, as well as a specific value and associated goal, and derive a metric that might be used to evaluate the recommender system relative to that goal. As previously noted, stakeholders are each assumed to have different values, corresponding value-driven goals and potential measures to reach these goals. It is worth reiterating that with these examples, we neither aim to provide a complete set of metrics that one might wish to implement in each of these settings nor highlight the most important metrics. Rather, we seek to illustrate the type of analysis needed to derive such metrics. Moreover, we expect the process of metric selection and development to be iterative rather than linear; this process may even take multiple rounds of consultation and implementation to derive a metric (or set of metrics) that captures a particular stakeholder’s perspective.
The three areas chosen are music streaming, educational resources, and job recommendations. These examples were chosen to highlight different stakeholder perspectives. In music streaming, we focus on musical artists as an example of the provider role. In the recommendation of educational materials, we have a domain where the value of a recommender is more than just consumer taste and yet personalization is still important; here we focus on the student/consumer. Job recommendation is, in many countries, subject to regulation intended to ensure non-discrimination in hiring and is therefore a good place to explore evaluation from the perspective of third-party stakeholders.

### v2

@@ 2.2 Societal and Human-Centric Values
Beyond economic and business values, societal and human-centric values, which cover other important aspects, are also crucial for businesses and platforms. Societal and human-centric values for stakeholders in recommender systems focus on ensuring that these systems operate in ways that prioritize humans individually and society as a whole. We find that there are 4 themes of societal and human-centric values for stakeholders in recommender systems that are relevant in the light of evaluation: (i) usefulness, (ii) well-being, (iii) legal and human rights, and (iv) public discourse and safety [109, 110, 28].
Usefulness and enjoyment means that recommendations should meet the needs and expectations of its stakeholders effectively and efficiently [61]. For example, in the case of a music recommender system, users should be able, via the recommender system, to discover new music that they might enjoy and match their tastes. At the same time, usefulness refers to the recommender system’s ability to help music artists get their outputs recommended to potentially interested listeners. Control and privacy are closely related values that pertains to the degree of influence and customization stakeholders might have over the recommendations that are generated. This includes privacy aspects in a way that users might want to control the amount of their (music) preference data that is shared with the recommender system [85, 86, 109].
Well-being refers to the extent to which a recommender system supports and enhances users’ psychological, emotional, and social health rather than simply optimizing engagement or consumption [28]. In the case of a music recommender system, this means that recommendations should influence the experience with the music streaming platform positively, e.g., provide music recommendations to help listeners relax or relieve stress [60]. In this respect, well-being is related to emotional, mental, psychological, physical, and social health. Other related values are connection, community and social bonding, e.g., to enable users to connect with like-minded music listeners or to enable music artists to contribute their outputs to a specific community. Thus, also reputation, recognition and acknowledgement might be valuable for some stakeholders, e.g., to support music artists in getting their contributions recognized by music listeners [81]. Personal growth and development might also be values contributing to well-being in the sense that, e.g., music recommendations could help people explore new music styles and genres, supporting exploration and self-discovery [16].
@@
Fairness can be related to diversity when the goal is to ensure, for example, that diverse articles and styles are represented in recommendation outputs. Diversity may also have listener-oriented benefits, e.g., help music listeners explore artists that might be new to them [93, 26]. A recommender system might seek to support freedom of expression as well as accessibility and inclusiveness by allowing, e.g., music artists to promote their content independent of the genre or popularity of their music [8, 94]. At the same time, recommender systems should enable users to access the content that they like and enjoy, even when their taste does not match the majority of music listeners [35, 65]. Transparency and trustworthiness might also be an important value for all stakeholders of a recommender system. For instance, music artists might be interested in why they are ranked at a specific position and music listeners might be interested in why a specific artist was recommended to them [84, 100].
Values in the area of public discourse and safety are related to a multitude of societal and human-centric aspects. Here, societal benefit goes beyond the satisfaction of individual stakeholders. As an example, a music streaming platform might be interested in fostering cultural enrichment by recommending a diverse set of music [113]. This is related to the value of tradition and history, for instance, by recommending local and traditional music, which might be hard to find without the recommender system [36, 71]. The environmental sustainability might also be an important value for some recommender systems stakeholders. This may involve implementing energy-efficient recommendation models within the platforms, or promoting local music artists whose concerts offer the opportunity for attendance without requiring extensive travel [77]. Finally, safety is concerned with users not being exposed to recommendations of disturbing ethically questionable, or age-inappropriate content. In the case of music recommendations, this could refer to music tracks with offensive lyrical content [78, 91].
@@ 2.3 Values in Practice
As we mentioned earlier, the concept of value can be perceived as abstract. We can think of values as high-level, somewhat theoretical, and qualitative objectives that, in practice, must be refined for quantification [33]. As mentioned by Beattie et al. [7], this process involves scoping “what exactly needs to be measured.” This inherently complex challenge becomes even more obvious in the context of multistakeholder recommender systems, where different stakeholders–such as producers and consumers–may prioritize distinct values, each requiring unique and/or conflicting evaluation approaches [7].
Nevertheless, when it comes to evaluation of multistakeholder recommender systems, we must find a way to quantify values. If the aim is to determine the ‘goodness’ for all involved; there is a need for ”methods for measuring how well a system is adhering to, promoting, or facilitating these values over time” [109]. Translating theoretical values into concrete, measurable concepts requires tailored quantitative and/or qualitative methodologies specific to each multistakeholder use case [33].
As an initial reference, we can look at the work by Wu et al. [115] focused on the fairness value. Although limited to consumer-sided and producer-sided fairness–two of the many stakeholders we mention in this narrative–their proposed framework for optimization and the metrics considered for overall evaluation, offer a starting point for exploration that could be extended to multiple stakeholders concurrently. The recent synthesis by Stray et al. [109], focused on human and societal-related values, provides valuable insights into how to translate these values into measurable metrics for assessment. However, much like we discuss earlier in this section, the authors emphasize that in a multistakeholder context, there is neither a definitive list of values nor a single method for quantifying them. Instead, they present possible interpretations of values within the context of (multistakeholder) recommender systems and suggest potential indicators for evaluating whether a system upholds or supports these values (for more details, see the Appendix in [109]).
Ultimately, there is no universal approach or single measure for evaluation, as the process depends on the stakeholders involved, the focus of the recommendation application, and many other factors. As an exhaustive checklist for holistically evaluating multistakeholder recommender systems is not feasible, we provide in Section 3 general guidance on navigating this evaluation process.
@@ 3 Methodology
As we previously noted, evaluating recommender systems is a contextually situated problem: different domains, recommendation tasks, and contexts require specific metrics and evaluation setups tailored to that specific recommendation scenario. Multistakeholder evaluation, where the perspectives of other stakeholders are taken into account in addition to that of the consumer, only increases the potential complexity of the evaluation. The complexity of multistakeholder evaluation is demonstrated by the richness and variety of the examples described in Section 4. As a result of this complexity, prescribing exactly which methods to use in which order is impractical. Instead, we attempt to describe best meta-practices for conducting successful multistakeholder evaluation in this section, divided into different stages. We consider this process to be iterative, as findings in a later stage can necessitate returning to an earlier stage, for instance, when learning of a new relevant stakeholder to include or when value shifts occur in one or more stakeholders. These stages are
  • Stakeholder identification
  • Value articulation
  • Operationalization of values as metrics
  • (if needed) Aggregation of metrics into an overall system objective
Recall that, in our discussion, we assume that we seek to evaluate an existing recommender system, one that has already been developed to provide a particular recommendation function. Of course, planning for a system’s evaluation should be part of its development: stakeholder consultation would naturally be prioritized in the design and implementation of a multistakeholder recommender system.
@@ 3.1 Stakeholders
Consumers (or users) have historically played (and will continue to play) a central role in recommender systems evaluation. As a result, a common next step would be profiling the consumer stakeholder and the different subgroups this stakeholder category may represent. In addition to interviews with the system stakeholders, any existing market or user research on the user base of the recommendation platform could serve as a valuable foundation for identifying representative subgroups within this user base. A literature review aimed at identifying similar or related recommendation scenarios could also help identify different user groups, especially groups that may be underrepresented in the market research for whatever reason. The system stakeholder should be able to facilitate access to these subgroups, for instance through user research panels, surveys on the website, or customer mailing lists. It is important to recruit a diverse and representative sample of consumers to represent the customer stakeholder and ensure all voices are heard in the evaluation process. Customers should be interviewed or surveyed about what values matter to them in this recommendation scenario (and their relative importance), what goals they have, and how and when they envision using the recommender system. If representative, the principle of saturation could be useful in guiding the sample size required: if additional participants do not reveal any new values, goals, or usage scenarios, then the sample should be representative of the customer stakeholder. Consumers are also a valuable source for identifying possible downstream stakeholders that are worth including in the evaluation process.
@@ 3.3 Evaluation Metrics
Formal evaluation metrics provide a way to measure the extent to which the goals of various stakeholders are achieved, i.e., they are measurable proxies towards goals. For example, both consumers and providers are likely to be interested in recommendation accuracy, consumers may be further interested in item discoverability (diversity, novelty, coverage); providers are likely interested in increasing revenue and engagement; and, the third-party stakeholders (for instance, regulators) are likely to be interested in protection-related metrics for consumers and providers (representation, fairness, etc.).
Multiple metrics can measure the success of the same goal, depending on the point of view or the aspect we want to highlight. For example, there are different metrics to measure accuracy, e.g., Normalized Discounted Cumulative Gain (NDCG), Mean Reciprocal Rank (MRR), or Recall; we may measure the overall number of items sold in a specific period or a specific geographical area, the items from the long-tail and the short-head, etc. Depending on the goal, we may have metrics subset of targeting the overall population of users and stakeholders available in the system.
@@
3.4 Strategies for Aggregate Evaluation
Identifying the list of important evaluation metrics $(m_{1},\ldots,m_{n})$, as discussed above, provides the ability to evaluate (i.e., to score) a given recommender system $R$ in a multidimensional manner. More formally, $\mathbf{S}(R)=(s_{1},\ldots,s_{n})$, where $s_{i}$ is the performance of $R$ with respect to measure $m_{i}$, i.e., $s_{i}=m_{i}(R)$. For some applications, this may be sufficient, as it provides a multidimensional view of the system’s performance.
In other cases, it may be desirable to formulate a single aggregate measure that captures the overall (i.e., multistakeholder, multiobjective) performance of the system [120]. In particular, given two candidate recommender systems $R_{A}$ and $R_{B}$, where each of which can be evaluated according to the stated list of metrics, $\mathbf{S}(R_{A})$ and $\mathbf{S}(R_{B})$, how to design a multistakeholder / multiobjective evaluation mechanism $\prec_{M}$ that allows to determine whether system $R_{B}$ has superior overall performance to system $R_{A}$, i.e., $\mathbf{S}(R_{A})\prec_{M}\mathbf{S}(R_{B})$? Designing multistakeholder / multiobjective evaluation mechanisms constitutes an important and non-trivial task, because the recommender systems (and ML-based predictive models, more generally) often exhibit significant tradeoffs when it comes to different performance metrics. To mention a few examples, prior literature has shown that there are tradeoffs between accuracy and item “discoverability” measures (i.e., diversity, serendipity, novelty) [4, 15, 75], between accuracy and fairness [20, 41, 76], between different fairness measures themselves [9, 59], and that business-oriented metrics may not be always aligned with the accuracy-oriented metrics [44, 51, 52].
From a high-level perspective, strategies for developing aggregate multistakeholder / multiobjective evaluation mechanisms $\prec_{M}$ can be classified into two general paradigms:
  • Direct/user-specified multistakeholder evaluation mechanisms, which typically rely on domain expertise to specify explicitly the relative importance and/or tradeoffs between individual performance metrics, which then can be used to formally define the overall multistakeholder evaluation mechanism $\prec_{M}$. Example strategies include:
@@ 3.4 Strategies for Aggregate Evaluation
  • Learning-based multistakeholder evaluation mechanisms, which typically infer the overall multistakeholder evaluation mechanism $\prec_{M}$ from certain available information. Example strategies include:
@@
Finally, once the multistakeholder evaluation mechanism is chosen/finalized, incorporating it into the overall recommender-system-based solution can be done in two general ways. (1) A straightforward post-processing approach would be to generate a number of different candidate recommender systems instances (e.g., using different recommendation algorithm configurations) and to use the multistakeholder evaluation mechanism to select the best-performing system among them. (2) In contrast, the algorithmic tuning approach would “push” the multistakeholder evaluation mechanism into the recommendation process, where a recommendation algorithm uses multistakeholder considerations more directly as part of the recommendation calculation/optimization process.
As an illustrative example, Figure 3 (adapted from [3, 51]) shows the performance of a simulated recommender system that takes into account the following multistakeholder considerations: item profitability (a key consideration for the recommendation provider) and recommendation relevance (a key consideration for the recommendation consumer). There is an inherent nuanced tradeoff between the two considerations. If the system is allowed to recommend only from the small number of the most relevant items (say, any items with predicted ratings above 4.9) for each user, the item relevance will be high, but choosing the most profitable items from only a small number of candidates will not improve the profitability metric. On the other extreme, if the system is allowed to recommend a much broader set of items (say, any items with predicted ratings above 3.0), the system will indeed be able to find more profitable items for recommendation, but the user may end up not choosing some of them due to their potentially low relevance (and may also be turned off by low-quality recommendations in the long run), thus, resulting in a lower actual profitability. The system uses the algorithmic tuning approach to determine the relevance constraint/threshold (i.e., 4.5) that maximizes the profit metric, taking into account both the individual item profitability and the fact that the users will reject inaccurate recommendations.
Figure 3: Recommendation relevance vs. profit (adapted from [3, 51])
@@ 3.5 Practical guidelines
  • Articulation of values and goals. The goals of the recommender system and expectations for what makes a recommendation good may vary considerably by stakeholder and context. Each stakeholder has different goals and expectations for the recommender system, and these are directly or indirectly tied to the values that matter to them. Qualitative research methods are particularly useful for identifying these values and goals. It is important to keep domain differences in mind when identifying the values and goals. In some domains, certain stakeholders might be affected more seriously by the selection of recommendations provided by the system. It is therefore recommended to undertake a risk analysis at this stage to properly understand the extent of these risks.
  • Selection of evaluation metrics. Once values and goals have been mapped, they have to be represented by measurable entities, e.g., by selecting existing metrics from prior research that are relevant for a given application context or by designing new ones. Part of this process includes determining which metrics to prioritize in the evaluation of the system, as typically multiple metrics can measure the success of the same goal depending on which perspective to highlight. These metrics can vary considerably in terms of being more qualitative or quantitative in nature, or in terms of representing more short-term or long-term interests. All selected metrics must be validated by the target stakeholders to check whether they are representative of their goals.
@@ 4 Example Applications and Metrics
Deriving an evaluation metric requires working from a construct, an abstract quality of the recommendation process that we would like to understand, to a concrete proxy of that construct that can be measured and designing a methodology to measure it. The application-specificity of multistakeholder evaluation makes it difficult to provide such analysis in a general way. With that in mind, we present several specific examples, which illustrate the design process for a variety of metrics for different stakeholder in different applications.
In these examples, we are not seeking to develop a full set of metrics for given application. As we note above, executing all of the stages of the evaluation design process can be a lengthy endeavor requiring detailed stakeholder consultation. For our hypothetical examples, we consider one thread of that much more involved process, focusing on a particular stakeholder, a specific value that they may have, and a particular metric that could be developed to evaluate it. It is worth reiterating that with these examples, we neither aim to provide a complete set of metrics that one might wish to implement in each of these settings nor to highlight the most important metrics. Rather, we seek to illustrate the type of analysis needed to derive such metrics. Moreover, we expect the process of metric selection and development to be iterative rather than linear; this process may take multiple rounds of consultation and implementation to derive a metric (or set of metrics) that captures a particular stakeholder’s perspective.
The three areas chosen are music streaming, educational resources, and job recommendations. These examples were chosen to highlight different stakeholder perspectives. In music streaming, we focus on musical artists as an example of the provider role. In the recommendation of educational materials, we have a domain where the value of a recommender is more than just consumer taste and yet personalization is still important; here we focus on the student, a consumer role. Job recommendation is, in many countries, subject to regulation intended to ensure non-discrimination in hiring and is therefore a good place to explore evaluation from the perspective of third-party stakeholders.